OpenStudy (anonymous):

PLZ HELP I WILL MEDAL AND FAN I HAVE TWO QUESTIONS PLZ! I WILL POST THEM NOW!

OpenStudy (anonymous):

What is the solution to the system of equations? 2x + y = -3 x - 2y = -4 (__,__)

pooja195 (pooja195):

x - 2y = -4 Start by isolating x add 2y to both sides what are you left with?

OpenStudy (anonymous):

i want and need to learn this, i just am confused here are just some random practice questions plz help thx

OpenStudy (anonymous):

i am left with x - 2y = -4 - 2y?

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

Ok wait lets start over OK So we need to isolate x so that we can solve using the substituion method :) Start by adding 2y to both sides so that we have x alone \[\huge~\rm~x - 2y +2y= -4+2y\]

OpenStudy (anonymous):

ok now what i am writing this down

pooja195 (pooja195):

Well what are we left with?

OpenStudy (anonymous):

x = 2y + 2y +2y + -4?

OpenStudy (anonymous):

oh i mean x = 6y + -4

OpenStudy (anonymous):

oops*

pooja195 (pooja195):

o-o not quite x=2y-4 We cancelled out the 2y

OpenStudy (anonymous):

oh oh oh, ok i thought i needed to add the 2y's up ok now what?

pooja195 (pooja195):

Now we have: Equation 1 \[\huge~\rm~x=2y-4\] Equation 2 \[\huge~\rm~2x + y = -3 \] Now we can substitute Equation 1 into equation 2 BECAUSE we know that x EQUALS 2y-4 so like this: \[\huge~\rm~2(2y-4)+y=-3\] Can you solve for y?

OpenStudy (anonymous):

yes one sec

OpenStudy (anonymous):

y=1?

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

Correct

OpenStudy (anonymous):

wait so what is x?

OpenStudy (anonymous):

(_?_, 1)

pooja195 (pooja195):

Thats the next part! :) Now we plug the y value into equation 1 =) \[\huge~\rm~x=2(1)−4\] solve for x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x= -2?

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

is it (-2,1)

pooja195 (pooja195):

Well done :)

OpenStudy (anonymous):

yay i have one more that needs checking

pooja195 (pooja195):

Ok :)